OpenStudy (anonymous):

Use the method of cylindrical shells to find the volume V of the solid obtained by rotating the region bounded by the given curves about the x-axis. x + y = 7, x = 16 − (y − 3)^2

OpenStudy (anonymous):

The two curves meet at : 6 - y = 15 - (y - 3)² y² - 7y = 0 y = 0 and y = 7 By the shells method , the radius is y , the height is 15 - (y - 3)² - (6 - y), the thickness is dy ; V = 2π ∫(y = 0 to 7) y ( 15 - (y - 3)² - (6 - y) ) dy = = 2π ∫(y = 0 to 7) 7y² - y³ dy = = 2π ( 7/3 y³ - 1/4 y⁴ ) | (y = 0 to 7) = = 2π ( 7/3 × 343 - 2401/4 ) = 2401π/6

OpenStudy (anonymous):

thanks.)